# Membership Privacy Risks of Sharpness Aware Minimization

### v3

@@ Appendix C Proofs
Under Lemma˜1, $\|\mathbf{x}_{3,i}\|^{2}=(1\pm o(1))\,d_{3}$ and the cross inner products are $o(\|\mathbf{x}_{3,i}\|\,\|\mathbf{x}_{3,j}\|)=o(d_{3})$
@@
Remark 1 (Distribution $F_{B}$ and why $F_{B}(A)$ appears).
For each minority sample $i$, the anti-alignment magnitude $B_{i}\triangleq-\,y_{i}\langle\mathbf{w}_{1},\bm{\nu}_{i}\rangle>0$ summarizes how strongly the majority anchor opposes the minority anchor for that sample. We assume $\{B_{i}\}_{i\in\mathcal{S}}$ are i.i.d. draws from a common distribution $F_{B}$ supported on $(0,B_{\max}]$. Unseen minority test point has margin $m^{\prime}=-B+A+\zeta^{\prime},\text{ with }B\sim F_{B},$ so $\Pr_{\mathcal{S}}(m^{\prime}>0)=\Pr(B<A)=F_{B}(A)$ (up to the $o(1)$ fluctuation $\zeta^{\prime}$ from Lemma˜1). Intuitively, $F_{B}(A)$ is the fraction of minority subclasses whose majority anti-alignment is not too strong relative to the learned shared minority signal $A$.
@@
Assumption 4 ($\beta$-stability under higher MSA).
Let $w^{(A)},w^{(B)}$ be interpolating solutions trained on the same $S$ with $A^{(A)}\geq A^{(B)}$. We assume that the per-point coefficient drift of the noise-block representer weights does not offset the threshold shift induced by $A$; namely, for every minority point $i\in\mathcal{S}$,
$\beta_{i}^{(A)}-\beta_{i}^{(B)}\;\geq\;-\;\frac{A^{(A)}-A^{(B)}}{\|x_{3,i}\|^{2}}\,.$ (8)
@@
Draw $(\mathbf{x}^{\prime},y^{\prime})\sim\mathcal{D}_{\mathcal{S}}$. $\mathbf{x}_{3}^{\prime}$ is independent of $\{\mathbf{x}_{3,j}\}$ and mean-zero; hence $y\langle\mathbf{w}_{3},\mathbf{x}_{3}^{\prime}\rangle=\sum_{j}\beta_{j}yy_{j}\langle\mathbf{x}_{3,j},\mathbf{x}_{3}^{\prime}\rangle$ is a mean-zero fluctuation with variance vanishing relative to $\|\mathbf{x}_{3}^{\prime}\|^{2}$ ; set this fluctuation to $\zeta^{\prime}=o_{\mathbb{P}}(1)$ using Lemma˜1. Then, $y^{\prime}\langle\mathbf{w},\mathbf{x}^{\prime}\rangle=-B+A+\zeta^{\prime}$.
@@
By Lemma˜3, minority test accuracy equals $F_{B}(A)$. This is monotone increasing in $A$. The majority subclass and noise samples yield the same result for both models by Lemmas˜2 and 4. ∎
Proof of Theorem˜2
Proof.
Decompose $\mathrm{Adv}^{\mathrm{M}}=\mathrm{TPR}-\mathrm{FPR}$.
TPR: On training points, majority subclass samples exceed any finite fixed $t$ by Assumption 2 and Lemma˜2.
For pure noise samples, fix any margin threshold $t\geq 0$. For a pure noise training point $i\in\mathcal{N}$,
$y_{i}\langle w,\mathbf{x}_{i}\rangle\;=\;\underbrace{y_{i}\langle\mathbf{w}_{1},\mathbf{x}_{1,i}\rangle+y_{i}\langle\mathbf{w}_{2},\mathbf{x}_{2,i}\rangle}_{G_{i}}\;+\;\underbrace{y_{i}\langle\mathbf{w}_{3},\mathbf{x}_{3,i}\rangle}_{=\ \beta_{i}\|\mathbf{x}_{3,i}\|^{2}+\zeta_{i}}\;=\;G_{i}\;+\;\beta_{i}\|\mathbf{x}_{3,i}\|^{2}\;+\;\zeta_{i}.$
Hence the threshold indicator admits the exact identity
$\mathds{1}\{\,y_{i}\langle\mathbf{w},\mathbf{x}_{i}\rangle\geq t\,\}\;=\;\mathds{1}\!\left\{\,\beta_{i}\;\geq\;\frac{t-G_{i}-\zeta_{i}}{\|\mathbf{x}_{3,i}\|^{2}}\right\}.$
By sub-Gaussianity of $\mathbf{x}_{1,i},\mathbf{x}_{2,i}$, there exists $r_{\eta}=\Theta(\|\mathbf{w}_{1}\|+\|\mathbf{w}_{2}\|)$ such that
$\Pr\!\big(\,|G_{i}|\leq r_{\eta}\,\big)\ \geq\ 1-\eta\qquad(\text{for any fixed }\eta\in(0,1)).$
By Lemma˜1, the remainder $\zeta_{i}=o_{\mathbb{P}}(1)$; that is, for every $\varepsilon>0$ there exists a function $\delta_{\varepsilon}=\delta_{\varepsilon}(d_{3})\downarrow 0$ as $d_{3}\to\infty$ such that
$\Pr\!\big(\,|\zeta_{i}|\leq\varepsilon\,\big)\ \geq\ 1-\delta_{\varepsilon}.$
Intersecting these events, with probability at least $1-\eta-\delta_{\varepsilon}$ we have
$t-r_{\eta}-\varepsilon\;\leq\;t-G_{i}-\zeta_{i}\;\leq\;t+r_{\eta}+\varepsilon.$
Plugging in yields
[…]
holding with probability $1-\eta-\delta_{\varepsilon}$. Letting $\varepsilon\downarrow 0$ and using $\|\mathbf{x}_{3,i}\|^{2}=(1\pm o(1))\,d_{3}$, define the data-dependent thresholds
$\tau_{i}^{(+)}(t)\;\coloneqq\;\frac{t+r_{\eta}}{\|\mathbf{x}_{3,i}\|^{2}},\qquad\tau_{i}^{(-)}(t)\;\coloneqq\;\frac{t-r_{\eta}}{\|\mathbf{x}_{3,i}\|^{2}},$
so that, with probability $1-\eta-o(1)$,
[…]
Therefore, the contribution of pure noise samples on TPR is not dependent on A. Under our setup, higher $A$ likely implies higher $\|\mathbf{w}_{2}\|$, increasing the required minimum for $\beta_{i}$. Higher $\beta_{i}$ can push the sample more above the threshold increasing TPR, more in line with our theorem. This effect, however, is much smaller than the effect from $A$ for minority subclass samples as it affects through $r_{\eta}=\Theta(\|\mathbf{w}_{1}\|+\|\mathbf{w}_{2}\|)$. Also by Assumption 4, the difference in A dominates.
For minority subclass samples,
$m_{i}\;=\;y_{i}\langle\mathbf{w},\mathbf{x}_{i}\rangle\;=\;-B_{i}+S+\beta_{i}\|\mathbf{x}_{3,i}\|^{2}+\zeta_{i}.$
Setting $m_{i}>t$ and rearranging, the total contribution of minority examples is
$\frac{1}{n_{\mathcal{S}}}\sum_{i\in\mathcal{S}}\mathds{1}\!\left\{\beta_{i}\geq\frac{t-(-B_{i}+A)}{\|x_{3,i}\|^{2}}\right\},$
which is pointwise nondecreasing in $A$. Hence $\mathrm{TPR}^{(A)}\geq\mathrm{TPR}^{(B)}$, with strict gain when at least one minority indicator flips. That is, if for $i\in\mathcal{S}$,
$\exists i\text{ }m_{i}^{(B)}<t<m_{i}^{(A)}$
FPR: For nonmembers, the mixture $\mathcal{D}$ splits into majority, minority, and noise. At a fixed $t$, the majority and noise tails do not depend on $A$. The minority nonmember term equals $F_{B}(A-t)+o(1)$ (Lemma˜3). If $t>\max\{A^{(A)},A^{(B)}\}$ and $B>0$ a.s., then $F_{B}(A^{(A)}\!-t)=F_{B}(A^{(B)}\!-t)=0$, so the minority part vanishes for both models. Therefore $\mathrm{FPR}^{(A)}=\mathrm{FPR}^{(B)}+o(1)$.
Combining, $\mathrm{Adv}^{\mathrm{M}}(\mathbf{w}^{(A)},t)\geq\mathrm{Adv}^{\mathrm{M}}(\mathbf{w}^{(B)},t)$, with strict inequality whenever some minority training indicator flips for $\mathbf{w}^{(A)}$ but not for $\mathbf{w}^{(B)}$. ∎

### v4

@@ Appendix C Proofs
Assumption 1.
We make the following assumptions:
  1. (i)
    (Kernel overlap) Let $P:=X^{\top}(XX^{\top})^{-1}X$ be the orthogonal projector onto $\mathrm{span}(X^{\top})$. The population-sharpness direction has nontrivial mass outside the data span:
    $\|(I-P)\,\Sigma\,\widehat{\theta}_{G_{0}}\|\ \geq\ c\ >\ 0$
    with probability 1o(1)1-o(1) for some constant $c$.
  2. (ii)
    (Bounded interpolator) The Euclidean interpolator satisfies $\|\widehat{\theta}_{G_{0}}\|=O_{\mathbb{P}}(1)$.
Remark 4 (Justification of assumptions).
Assumption 1(i) is generic in the overparameterized regime ($d\gg n$): the vector $\Sigma\widehat{\theta}_{G_{0}}$ is a population signal direction, while $\mathrm{span}(X^{\top})$ is a random $n$-dimensional subspace. Under mild spectral regularity of $\Sigma$, their alignment is not perfect with high probability, yielding a non-negligible (IP)(I-P) component.
Assumption 1(ii) is standard for benign overfitting and ensures finite test variance. If $\|\widehat{\theta}_{G_{0}}\|$ diverged, the Euclidean interpolator would have exploding prediction variance and thus be unstable on unseen data.
Lemma 1.
The unique solution of equation 6 is
$\widehat{\theta}_{G}=G^{-1}X^{\top}\Bigl(XG^{-1}X^{\top}\Bigr)^{-1}y,\qquad\text{and}\qquad X\widehat{\theta}_{G}=y.$
Proof.
Form the Lagrangian $L(\theta,\lambda)=\tfrac{1}{2}\theta^{\top}G\theta+\lambda^{\top}(X\theta-y)$. The KKT conditions are $G\theta+X^{\top}\lambda=0$ and $X\theta=y$. Eliminate $\theta=-G^{-1}X^{\top}\lambda$ to get $XG^{-1}X^{\top}\lambda=y$, hence $\lambda=-(XG^{-1}X^{\top})^{-1}y$. Substituting back yields the stated $\widehat{\theta}_{G}$. Since $G\succ 0$, the objective is strictly convex and the solution is unique. ∎
Lemma 2.
Let $\widehat{\theta}(\eta):=\widehat{\theta}_{G_{\eta}}$ with $G_{\eta}=I+\eta\Sigma$. Then
$\widehat{\theta}^{\prime}(0):=\frac{d}{d\eta}\widehat{\theta}(\eta)\Big|_{\eta=0}=-(I-P)\,\Sigma\,\widehat{\theta}_{G_{0}}.$
Proof.
By Lemma 1,
$\widehat{\theta}(\eta)=G_{\eta}^{-1}X^{\top}\!\bigl(XG_{\eta}^{-1}X^{\top}\bigr)^{-1}y.$
Differentiate at $\eta=0$.
(i) $\frac{d}{d\eta}G_{\eta}^{-1}=-G_{\eta}^{-1}\Sigma G_{\eta}^{-1}$, so $\bigl.\frac{d}{d\eta}G_{\eta}^{-1}\bigr|_{\eta=0}=-\Sigma$.
(ii) Let $A(\eta):=XG_{\eta}^{-1}X^{\top}$. Then $A^{\prime}(0)=-X\Sigma X^{\top}$ and
$\bigl.\tfrac{d}{d\eta}A(\eta)^{-1}\bigr|_{\eta=0}=A(0)^{-1}(X\Sigma X^{\top})A(0)^{-1}.$
Combining (i)–(ii),
$\widehat{\theta}^{\prime}(0)$ $=-\Sigma\,X^{\top}(XX^{\top})^{-1}y+X^{\top}(XX^{\top})^{-1}(X\Sigma X^{\top})(XX^{\top})^{-1}y$
$=-(I-P)\Sigma\,\widehat{\theta}_{G_{0}}.$
Lemma 3 (Distribution of model output on non-member data).
Let $X_{\mathrm{out}}\sim\mathcal{N}(0,\Sigma)$ independently of the training set. Conditioned on $\widehat{\theta}_{G}$, the prediction $f_{G}(X_{\mathrm{out}})=\widehat{\theta}_{G}^{\top}X_{\mathrm{out}}$ satisfies
$f_{G}(X_{\mathrm{out}})\sim\mathcal{N}\!\bigl(0,\sigma_{G}^{2}\bigr),\qquad\sigma_{G}^{2}:=\widehat{\theta}_{G}^{\top}\Sigma\,\widehat{\theta}_{G}.$
Proof.
Condition on $\widehat{\theta}_{G}$. Since $X_{\mathrm{out}}$ is Gaussian and $f_{G}$ is linear in $X_{\mathrm{out}}$, the claim follows with variance $\widehat{\theta}_{G}^{\top}\Sigma\widehat{\theta}_{G}$. ∎
Lemma 4 (First derivative of non-member variance at $\eta=0$).
With $G_{\eta}=I+\eta\Sigma$, let $\sigma^{2}(\eta):=\widehat{\theta}(\eta)^{\top}\Sigma\,\widehat{\theta}(\eta)$. Then
$\sigma^{2\,\prime}(0)=2\,\widehat{\theta}_{G_{0}}^{\top}\Sigma\,\widehat{\theta}^{\prime}(0)=-2\,\big\|(I-P)\Sigma\,\widehat{\theta}_{G_{0}}\big\|^{2}.$
Proof.
Differentiate: $\sigma^{2\,\prime}(0)=2\,\widehat{\theta}_{G_{0}}^{\top}\Sigma\,\widehat{\theta}^{\prime}(0)$ since $\Sigma$ is symmetric. Insert Lemma 2:
$\sigma^{2\,\prime}(0)=-2\,\widehat{\theta}_{G_{0}}^{\top}\Sigma(I-P)\Sigma\,\widehat{\theta}_{G_{0}}=-2\|(I-P)\Sigma\widehat{\theta}_{G_{0}}\|^{2},$
using symmetry and idempotence of (IP)(I-P). ∎
Proof of Theorem 1
Proof.
By Lemma 4,
$\sigma^{2\,\prime}(0)=-2\|(I-P)\Sigma\widehat{\theta}_{G_{0}}\|^{2}.$
Assumption 1(i) yields $\|(I-P)\Sigma\widehat{\theta}_{G_{0}}\|\geq c$, hence $\sigma^{2\,\prime}(0)\leq-2c^{2}<0$ with probability 1o(1)1-o(1).
Since $\widehat{\theta}(\eta)$ is smooth in $\eta$ for $G_{\eta}\succ 0$, $\sigma^{2}(\eta)$ is differentiable at 0. Therefore, on the same high-probability event, there exists $\eta_{0}>0$ such that for all $\eta\in(0,\eta_{0}]$,
$\sigma^{2}(\eta)=\sigma^{2}(0)+\eta\,\sigma^{2\,\prime}(0)+o(\eta)<\sigma^{2}(0).$
This implies $\sigma_{G_{\eta}}^{2}<\sigma_{G_{0}}^{2}$ for all sufficiently small $\eta>0$ with probability 1o(1)1-o(1). ∎
C.1 Confidence-threshold attack
Let $\mathcal{I}_{\mathrm{in}}$ be the index set of member training points, and let $\{(x_{i},y_{i})\}_{i\in\mathcal{I}_{\mathrm{in}}}$ denote the training samples. For a geometry $G\succ 0$, let $\widehat{\theta}_{G}\in\mathbb{R}^{d}$ be the (interpolating) solution of equation 6. Define the signed score and confidence by
$f_{G}(x):=\widehat{\theta}_{G}^{\top}x,\qquad\mathrm{Conf}_{G}(x):=\bigl|f_{G}(x)\bigr|=\bigl|\widehat{\theta}_{G}^{\top}x\bigr|.$
We model a black-box confidence-threshold attacker as follows. Let $I$ be a random index drawn from $\mathcal{I}_{\mathrm{in}}$ (e.g., uniformly), and let the random member pair be
$(X_{\mathrm{in}},Y_{\mathrm{in}}):=(x_{I},y_{I}).$
Let $X_{\mathrm{out}}\sim\mathcal{N}(0,\Sigma)$ be an independent non-member (test) input, independent of the training set and algorithmic randomness. For a threshold $\tau\geq 0$, define
$\mathrm{TPR}_{G}(\tau)$ $:=\Pr\!\left(\mathrm{Conf}_{G}(X_{\mathrm{in}})\geq\tau\right),$ (7)
$\qquad\mathrm{FPR}_{G}(\tau)$ $:=\Pr\!\left(\mathrm{Conf}_{G}(X_{\mathrm{out}})\geq\tau\right),$ (8)
$\qquad\mathrm{Adv}^{\mathrm{conf}}_{G}$ $:=\sup_{\tau\geq 0}\Bigl(\mathrm{TPR}_{G}(\tau)-\mathrm{FPR}_{G}(\tau)\Bigr).$ (9)
Lemma 5.
For any geometry $G\succ 0$,
$\mathrm{Conf}_{G}(X_{\mathrm{in}})=|Y_{\mathrm{in}}|\quad\text{almost surely}.$
Proof.
Because $\widehat{\theta}_{G}$ interpolates the training data, we have $\widehat{\theta}_{G}^{\top}x_{i}=y_{i}$ for every $i\in\mathcal{I}_{\mathrm{in}}$. Therefore for the random member index $I$,
$\mathrm{Conf}_{G}(X_{\mathrm{in}})=\bigl|\widehat{\theta}_{G}^{\top}x_{I}\bigr|=|y_{I}|=|Y_{\mathrm{in}}|\quad\text{a.s}.$
Lemma 6.
Conditioned on $\widehat{\theta}_{G}$,
$\mathrm{Conf}_{G}(X_{\mathrm{out}})\overset{d}{=}|Z|,\ \ Z\sim\mathcal{N}(0,\sigma_{G}^{2}),$
Proof.
By Lemma 3 and absolute value. ∎
Lemma 7.
Let $0<\sigma_{1}<\sigma_{2}$ and let $Z_{k}\sim\mathcal{N}(0,\sigma_{k}^{2})$ for $k\in\{1,2\}$. Then for every $\tau>0$,
$\Pr(|Z_{1}|\geq\tau)<\Pr(|Z_{2}|\geq\tau).$
Proof.
Let $U\sim\mathcal{N}(0,1)$. Then $Z_{k}\overset{d}{=}\sigma_{k}U$, so
$\Pr(|Z_{k}|\geq\tau)=\Pr\!\left(|U|\geq\frac{\tau}{\sigma_{k}}\right).$
The function $t\mapsto\Pr(|U|\geq t)$ is strictly decreasing on $(0,\infty)$. Since $\tau/\sigma_{1}>\tau/\sigma_{2}$, the claim follows. ∎
Lemma 8.
Assume $\Pr(|Y_{\mathrm{in}}|>0)>0$ (true whenever labels have any continuous noise). Then
$\mathrm{Adv}^{\mathrm{conf}}_{G}=\sup_{\tau>0}\Bigl(\mathrm{TPR}_{G}(\tau)-\mathrm{FPR}_{G}(\tau)\Bigr).$
Proof.
At $\tau=0$, $\mathrm{TPR}_{G}(0)=\mathrm{FPR}_{G}(0)=1$, so the gap equals 0. Because $\Pr(|Y_{\mathrm{in}}|>0)>0$, we have $\mathrm{TPR}_{G}(\tau)>0$ for some $\tau>0$. Also, by Lemma 6, $\mathrm{FPR}_{G}(\tau)\to 0$ as $\tau\to\infty$. Hence there exists a $\tau>0$ such that $\mathrm{TPR}_{G}(\tau)-\mathrm{FPR}_{G}(\tau)>0$. Therefore the supremum cannot be attained at $\tau=0$, and we may restrict to $\tau>0$. ∎
Proof of Theorem 2
Proof.
By Lemma 5, member confidences are geometry-invariant, so
$\mathrm{TPR}_{G_{\eta}}(\tau)=\mathrm{TPR}_{G_{0}}(\tau)\quad\text{for all }\tau\geq 0.$
By Theorem 1, for all sufficiently small $\eta>0$,
$\sigma_{G_{\eta}}^{2}<\sigma_{G_{0}}^{2}\quad\text{with probability }1-o(1).$
Condition on this high-probability event. Lemma 6 implies $\mathrm{Conf}_{G_{\eta}}(X_{\mathrm{out}})\overset{d}{=}|Z_{\eta}|$ with $Z_{\eta}\sim\mathcal{N}(0,\sigma_{G_{\eta}}^{2})$, and similarly for $G_{0}$. Then Lemma 7 yields that for every $\tau>0$,
$\mathrm{FPR}_{G_{\eta}}(\tau)=\Pr(|Z_{\eta}|\geq\tau)<\Pr(|Z_{0}|\geq\tau)=\mathrm{FPR}_{G_{0}}(\tau).$
Therefore, for every $\tau>0$,
$\mathrm{TPR}_{G_{\eta}}(\tau)-\mathrm{FPR}_{G_{\eta}}(\tau)>\mathrm{TPR}_{G_{0}}(\tau)-\mathrm{FPR}_{G_{0}}(\tau).$
Taking the supremum over $\tau>0$ and using Lemma 8,
$\mathrm{Adv}^{\mathrm{conf}}_{G_{\eta}}=\sup_{\tau>0}\Bigl(\mathrm{TPR}_{G_{\eta}}(\tau)-\mathrm{FPR}_{G_{\eta}}(\tau)\Bigr)>\sup_{\tau>0}\Bigl(\mathrm{TPR}_{G_{0}}(\tau)-\mathrm{FPR}_{G_{0}}(\tau)\Bigr)=\mathrm{Adv}^{\mathrm{conf}}_{G_{0}}.$
This holds with probability 1o(1)1-o(1). ∎
C.2 Likelihood-ratio attack
We consider an oracle likelihood-ratio attack that uses the model output $f_{G}(x)=\widehat{\theta}_{G}^{\top}x$. The attacker knows the true member and non-member score distributions and performs the Neyman–Pearson likelihood-ratio test.
Score distributions.
Recall the data model $x\sim\mathcal{N}(0,\Sigma)$ and $y=\theta^{*\top}x+\xi$ with $\xi\sim\mathcal{N}(0,\sigma_{y}^{2})$ independent of $x$. Let $(X_{\mathrm{in}},Y_{\mathrm{in}})$ be a random member training pair obtained by sampling a random index from the training set. Unconditionally over the training data draw, we have the same marginal law as a fresh sample:
$X_{\mathrm{in}}\sim\mathcal{N}(0,\Sigma),\qquad Y_{\mathrm{in}}=\theta^{*\top}X_{\mathrm{in}}+\xi.$
Lemma 9 (Member score is geometry-invariant Gaussian).
For any geometry $G$, the member score satisfies
$f_{G}(X_{\mathrm{in}})=Y_{\mathrm{in}}\quad\text{a.s.}\qquad\text{and hence}\qquad f_{G}(X_{\mathrm{in}})\sim\mathcal{N}(0,v_{\mathrm{in}}),$
where $v_{\mathrm{in}}:=\theta^{*\top}\Sigma\theta^{*}+\sigma_{y}^{2}.$
Proof.
Interpolation gives $X\widehat{\theta}_{G}=y$, so for every training point $\widehat{\theta}_{G}^{\top}x_{i}=y_{i}$. Thus for a random member index $I$,
$f_{G}(X_{\mathrm{in}})=\widehat{\theta}_{G}^{\top}x_{I}=y_{I}=Y_{\mathrm{in}}\quad\text{a.s.}$
Unconditionally, $Y_{\mathrm{in}}=\theta^{*\top}X_{\mathrm{in}}+\xi$ with $X_{\mathrm{in}}\sim\mathcal{N}(0,\Sigma)$ and $\xi\sim\mathcal{N}(0,\sigma_{y}^{2})$ independent, so it is mean-zero Gaussian with variance $v_{\mathrm{in}}$. ∎
Lemma 10 (Non-member score is geometry-dependent Gaussian).
Let $X_{\mathrm{out}}\sim\mathcal{N}(0,\Sigma)$ be independent of the training set. Conditioned on $\widehat{\theta}_{G}$,
$f_{G}(X_{\mathrm{out}})=\widehat{\theta}_{G}^{\top}X_{\mathrm{out}}\sim\mathcal{N}(0,v_{\mathrm{out}}(G)),\qquad v_{\mathrm{out}}(G):=\widehat{\theta}_{G}^{\top}\Sigma\widehat{\theta}_{G}.$
Proof.
Same argument as Lemma 3. ∎
Oracle likelihood-ratio test.
Let $p_{\mathrm{in}}$ and $p_{\mathrm{out}}$ be the densities of $\mathcal{N}(0,v_{\mathrm{in}})$ and $\mathcal{N}(0,v_{\mathrm{out}}(G))$, respectively. The oracle LR score is
$\Lambda_{G}(s):=\log\frac{p_{\mathrm{in}}(s)}{p_{\mathrm{out}}(s)}.$
Lemma 11.
For zero-mean Gaussians with variances $v_{\mathrm{in}}>0$ and $v_{\mathrm{out}}(G)>0$, the LR test $\Lambda_{G}(s)\geq t$ is equivalent to $|s|\geq\tau$ for some $\tau\geq 0$. Moreover, the optimal test for any fixed FPR is of this form.
Proof.
For $s\in\mathbb{R}$,
$\Lambda_{G}(s)=-\frac{1}{2}\log v_{\mathrm{in}}+\frac{1}{2}\log v_{\mathrm{out}}(G)-\frac{s^{2}}{2v_{\mathrm{in}}}+\frac{s^{2}}{2v_{\mathrm{out}}(G)}=C_{G}+\frac{s^{2}}{2}\Bigl(\frac{1}{v_{\mathrm{out}}(G)}-\frac{1}{v_{\mathrm{in}}}\Bigr),$
where $C_{G}$ does not depend on $s$. Thus $\Lambda_{G}(s)\geq t$ is equivalent to $s^{2}\geq\tau^{2}$ for some $\tau\geq 0$, i.e. $|s|\geq\tau$. Neyman–Pearson gives optimality of the LR test. Note that as we are in a setting where members are highly confident, $v_{\mathrm{in}}>v_{\mathrm{out}}$. ∎
Define the LR-attack advantage as the best achievable TPR–FPR gap over all two-sided thresholds:
$\mathrm{Adv}^{\mathrm{LR}}_{G}:=\sup_{\tau\geq 0}\Bigl(\Pr(|f_{G}(X_{\mathrm{in}})|\geq\tau)-\Pr(|f_{G}(X_{\mathrm{out}})|\geq\tau)\Bigr).$
Proof of Theorem 3
Proof.
By Lemma 9,
$f_{G_{\eta}}(X_{\mathrm{in}})\sim\mathcal{N}(0,v_{\mathrm{in}})\quad\text{for all }\eta\geq 0,$
so the member tail $\Pr(|f_{G_{\eta}}(X_{\mathrm{in}})|\geq\tau)$ is geometry-invariant for every $\tau\geq 0$.
By Lemma 10,
$f_{G_{\eta}}(X_{\mathrm{out}})\sim\mathcal{N}(0,v_{\mathrm{out}}(G_{\eta})),\qquad v_{\mathrm{out}}(G_{\eta})=\sigma_{G_{\eta}}^{2}.$
Theorem 1 gives that, for all sufficiently small $\eta>0$,
$v_{\mathrm{out}}(G_{\eta})<v_{\mathrm{out}}(G_{0})\quad\text{with probability }1-o(1).$
Condition on this high-probability event. Then Lemma 7 (applied to $|f_{G}(X_{\mathrm{out}})|$) yields that for every $\tau>0$,
$\Pr(|f_{G_{\eta}}(X_{\mathrm{out}})|\geq\tau)<\Pr(|f_{G_{0}}(X_{\mathrm{out}})|\geq\tau).$
Therefore, for every $\tau>0$,
$\Pr(|f_{G_{\eta}}(X_{\mathrm{in}})|\geq\tau)-\Pr(|f_{G_{\eta}}(X_{\mathrm{out}})|\geq\tau)>\Pr(|f_{G_{0}}(X_{\mathrm{in}})|\geq\tau)-\Pr(|f_{G_{0}}(X_{\mathrm{out}})|\geq\tau).$
Taking the supremum over $\tau>0$ on both sides gives $\mathrm{Adv}^{\mathrm{LR}}_{G_{\eta}}>\mathrm{Adv}^{\mathrm{LR}}_{G_{0}}.$ (As in Lemma 8, $\tau=0$ yields zero gap, so the supremum is attained for some $\tau>0$.) This holds with probability 1o(1)1-o(1). ∎
Corollary 5 (Sample-adaptive LR monotonicity).
Fix a query point $z$:
$f_{\mathrm{in},z}\sim\mathcal{N}(0,v_{\mathrm{in},z}),\qquad f_{\mathrm{out},z}(G)\sim\mathcal{N}(0,v_{\mathrm{out},z}(G)).$
Let $G_{a},G_{b}$ be two geometries such that $v_{\mathrm{out},z}(G_{a})<v_{\mathrm{out},z}(G_{b})$. Then for every false positive rate $\alpha\in(0,1)$, the optimal likelihood-ratio test at level $\alpha$ attains strictly larger true positive rate under $G_{a}$ than under $G_{b}$. Equivalently, the per-sample ROC curve under $G_{a}$ strictly dominates that under $G_{b}$, so any MIA advantage is strictly larger under $G_{a}$.
Proof.
Fix $z$ and $k\in\{a,b\}$. The LR between $\mathcal{N}(0,v_{\mathrm{in},z})$ and $\mathcal{N}(0,v_{\mathrm{out},z}(G_{k}))$ is
$\Lambda_{z}(s;G_{k})=\frac{1}{2}\log\frac{v_{\mathrm{out},z}(G_{k})}{v_{\mathrm{in},z}}+\frac{s^{2}}{2}\Bigl(\frac{1}{v_{\mathrm{out},z}(G_{k})}-\frac{1}{v_{\mathrm{in},z}}\Bigr).$
If $v_{\mathrm{in},z}\neq v_{\mathrm{out},z}(G_{k})$ then $\Lambda_{z}(s;G_{k})$ is a strictly monotone function of $|s|$, so by the Neyman–Pearson lemma the optimal level-$\alpha$ test is equivalent to a two-sided magnitude test $|s|\geq\tau_{k}(\alpha)$ for some unique threshold $\tau_{k}(\alpha)>0$.
Write $f_{\mathrm{out},z}(G_{k})\overset{d}{=}\sqrt{v_{\mathrm{out},z}(G_{k})}\,U$ with $U\sim\mathcal{N}(0,1)$. The constraint $\Pr(|f_{\mathrm{out},z}(G_{k})|\geq\tau_{k}(\alpha))=\alpha$ is then
$\alpha=\Pr\Bigl(|U|\geq\tfrac{\tau_{k}(\alpha)}{\sqrt{v_{\mathrm{out},z}(G_{k})}}\Bigr).$
Since $u\mapsto\Pr(|U|\geq u)$ is strictly decreasing on $(0,\infty)$ and $v_{\mathrm{out},z}(G_{a})<v_{\mathrm{out},z}(G_{b})$, this forces
$\frac{\tau_{a}(\alpha)}{\sqrt{v_{\mathrm{out},z}(G_{a})}}=\frac{\tau_{b}(\alpha)}{\sqrt{v_{\mathrm{out},z}(G_{b})}}\quad\Rightarrow\quad\tau_{a}(\alpha)<\tau_{b}(\alpha).$
The member distribution $f_{\mathrm{in},z}$ is the same under $G_{a}$ and $G_{b}$, so
$\Pr\bigl(|f_{\mathrm{in},z}|\geq\tau_{a}(\alpha)\bigr)>\Pr\bigl(|f_{\mathrm{in},z}|\geq\tau_{b}(\alpha)\bigr).$
Thus at every FPR level $\alpha$ the optimal LR/LiRA test has strictly larger TPR under $G_{a}$, which implies strict ROC dominance and the claimed advantage comparison. ∎
C.3 Higher Subclass Alignment leads to higher generalization
@@
Under Lemma˜12, $\|\mathbf{x}_{3,i}\|^{2}=(1\pm o(1))\,d_{3}$ and the cross inner products are $o(\|\mathbf{x}_{3,i}\|\,\|\mathbf{x}_{3,j}\|)=o(d_{3})$
@@
Remark 5.
For each minority sample $i$, the anti-alignment magnitude $B_{i}\triangleq-\,y_{i}\langle\mathbf{w}_{1},\bm{\nu}_{i}\rangle>0$ summarizes how strongly the majority anchor opposes the minority anchor for that sample. We assume $\{B_{i}\}_{i\in\mathcal{S}}$ are i.i.d. draws from a common distribution $F_{B}$ supported on $(0,B_{\max}]$. Unseen minority test point has margin $m^{\prime}=-B+A+\zeta^{\prime},\text{ with }B\sim F_{B},$ so $\Pr_{\mathcal{S}}(m^{\prime}>0)=\Pr(B<A)=F_{B}(A)$ (up to the $o(1)$ fluctuation $\zeta^{\prime}$ from Lemma˜12). Intuitively, $F_{B}(A)$ is the fraction of minority subclasses whose majority anti-alignment is not too strong relative to the learned shared minority signal $A$.
@@
Draw $(\mathbf{x}^{\prime},y^{\prime})\sim\mathcal{D}_{\mathcal{S}}$. $\mathbf{x}_{3}^{\prime}$ is independent of $\{\mathbf{x}_{3,j}\}$ and mean-zero; hence $y\langle\mathbf{w}_{3},\mathbf{x}_{3}^{\prime}\rangle=\sum_{j}\beta_{j}yy_{j}\langle\mathbf{x}_{3,j},\mathbf{x}_{3}^{\prime}\rangle$ is a mean-zero fluctuation with variance vanishing relative to $\|\mathbf{x}_{3}^{\prime}\|^{2}$ ; set this fluctuation to $\zeta^{\prime}=o_{\mathbb{P}}(1)$ using Lemma˜12. Then, $y^{\prime}\langle\mathbf{w},\mathbf{x}^{\prime}\rangle=-B+A+\zeta^{\prime}$.
@@
By Lemma˜14, minority test accuracy equals $F_{B}(A)$. This is monotone increasing in $A$. The majority subclass and noise samples yield the same result for both models by Lemmas˜13 and 15. ∎